### v3

@@ 2.1 Network Architecture and Topology Modification
Neural populations were modeled as undirected stochastic block model (SBM) networks [5] with two equally sized clusters ($N=500$ is the total number of neurons, $|V_{t}|=|V_{o}|=250$, where $|V_{t}|$ and $|V_{o}|$ are the sizes of the stimulated $V_{t}$ and measured $V_{o}$ populations of neurons). The adjacency matrix $A$ was constructed with intra-cluster connection probability $p_{\mathrm{intra}}=0.15$ and inter-cluster probability $p_{\mathrm{inter}}$ systematically varied from 0.01 to 0.1. The signal was applied to neurons within the top 10%, 15%, or 20% centrality percentiles ($p_{\mathrm{input}}$) or to 10–20% of randomly selected neurons within the cluster $V_{t}$ (Proxy stimulation strategy and Random baseline). To overcome SBM’s inherent homogeneity and in order to introduce topological diversity, we implemented a centrality boosting procedure defined by:
@@
  • $\beta=1.5$ (50% centrality boost)
  • $c(u)$ is centrality value of the $u$ vertex
  • $S$ is the set of selected driver neurons
  • $k_{\mathrm{ext}}(u)$ is the existing inter-cluster degree
@@ 2.2 Centrality Measure Computation
[…]
@@ 2.3 Neural Dynamics Model
$v_{i}$$\mathrel{-}=\sum_{i}w_{i}\,\delta\bigl{(}t-t_{i}\bigr{)}.$ (11)
[…]
@@
In addition to noise from external influences, the driver neurons received an additional periodic signal:
@@
[…]
@@ 2.4 Numerical Experiments
We conducted several experiments in which we chose specific driver nodes in the population $V_{t}$ (with a fraction $p_{input}$ of the total amount) and measured neural activity in the population $V_{o}$. Several experimental protocols were implemented. For the first set of experiments, the SBM network was generated with the specified $p_{\mathrm{intra}}$, $p_{\mathrm{inter}}$. Next, the centrality measures were computed for the nodes of the $V_{t}$ population. 10-to-20% of the nodes with the highest centralities in the population $V_{t}$ were chosen as driver nodes. The dynamics of the resulting network was simulated for 5 seconds of biological time in Brian 2 neural simulator [9]. The initial 100 ms warm-up period was excluded from the analysis.
@@ 3 Results
    Threshold modulation: Boosted networks exhibited significant propagation amplification, compared to non-boosted case 3.15$\times$ increase in propagation efficiency compared to non-boosted networks at $p_{\mathrm{inter}}=0.10$, with the mean firing rate in $V_{o}$ rising from $3.33\pm 0.98$ Hz to $10.5\pm 0.13$ Hz under optimal driver selection. This demonstrates that targeted neuroplasticity can reconfigure information routing pathways to overcome physical connectivity constraints.
@@ 5 Conclusion
  • Critical operating regime: Centrality-based selection achieves maximum advantage (64×64$\times$64 × efficiency gain over random baseline, $p_{\mathrm{value}}<10^{-9}$) near network percolation thresholds
  • Metric-specific performance: Closeness and harmonic centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement (>5×>5$\times$> 5 × advantage over proxy strategies)
  • Mechanistic distinction: Centrality-based stimulation preserves temporal signal fidelity while proxy strategies amplify broadband noise

### v4

@@ 2.1 Network Architecture and Topology Modification
Neural populations were modeled as undirected stochastic block model (SBM) networks [5] with two equally sized clusters ($N=500$ is the total number of neurons, $|V_{t}|=|V_{o}|=250$, where $|V_{t}|$ and $|V_{o}|$ are the sizes of the stimulated $V_{t}$ and measured $V_{o}$ populations of neurons). The adjacency matrix $A$ was constructed with intra-cluster connection probability $p_{\mathrm{intra}}=0.15$ and inter-cluster probability $p_{\mathrm{inter}}$ systematically varied from 0.01 to 0.1. The signal was applied to neurons within the top 10%, 15%, or 20% centrality percentiles ($p_{\mathrm{input}}$) or to 10–20% of randomly selected neurons within the cluster $V_{t}$. To overcome SBM’s inherent homogeneity and in order to introduce topological diversity, we implemented a centrality boosting procedure defined by:
@@
  • $\beta=1.5$ (50% centrality boost);
  • $c(u)$ is centrality value of the $u$ vertex;
  • $S$ is the set of selected driver neurons;
  • $k_{\mathrm{ext}}(u)$ is the existing inter-cluster degree.
@@ 2.2 Centrality Measure Computation
$C_{c}(u)=\frac{n-1}{\sum_{v\neq u}d(u,v)}.\quad\text{(undefined for disconnected components)}$ (5)
@@ 2.3 Neural Dynamics Model
$v_{i}$ -$\mathrel{-}=\sum_{i}w_{i}\,\delta\bigl{(}t-t_{i}\bigr{)}.$ (11)
To simulate background neural activity, all neurons received independent stimuli $v_{n}$ with a Poisson distribution. Background activity was modeled as Poisson input with rate $\lambda_{\mathrm{bg}}=$1\text{\,}\mathrm{Hz}$\$ per neuron. Each spike from the Poisson generator causes an instantaneous increase in membrane potential, which can be represented as the sum of delta functions:
@@
In addition to the noise from external influences, the driver neurons received an additional periodic signal:
@@
Membrane capacitance $C_{m}$ $250\text{\,}\mathrm{pF}$
Membrane resistance $R_{m}$ $80\text{\,}\mathrm{M\SIUnitSymbolOhm}$
Resting potential $v_{rest}$ $65\text{\,}\mathrm{mV}$
Threshold $v_{th}$ $50\text{\,}\mathrm{mV}$
Reset potential $v_{reset}$ $70\text{\,}\mathrm{mV}$
Membrane time constant $\tau_{m}$ $20\text{\,}\mathrm{ms}$
Excit. and inhib. weights $w$ $1\text{\,}$
Input current $I_{0}$ $1\text{\,}\mathrm{pA}$
frequency $f$ $10\text{\,}\mathrm{Hz}$
phase $\phi$ $0\text{\,}$
@@ 2.4 Numerical Experiments
We conducted several experiments in which we chose specific driver nodes in the population $V_{t}$ (with a fraction $p_{input}$ of the total amount) and measured neural activity in the population $V_{o}$. Several experimental protocols were implemented. For the first set of experiments, the SBM network was generated with the specified probabilities $p_{\mathrm{intra}}$, $p_{\mathrm{inter}}$. Next, the centrality measures were computed for the nodes of the $V_{t}$ population. 10-to-20% of the nodes with the highest centralities in the population $V_{t}$ were chosen as driver nodes. The dynamics of the resulting network was simulated for 5 seconds of biological time in Brian 2 neural simulator [9]. The initial 100 ms warm-up period was excluded from the analysis.
@@ 3 Results
[…]
@@ 5 Conclusion
  • Critical operating regime: Centrality-based selection achieves maximum advantage ($64\times$ efficiency gain over random baseline, $p_{\mathrm{value}}<10^{-9}$) near network percolation thresholds;
  • Metric-specific performance: Closeness and harmonic centrality consistently outperformed other metrics, with closeness centrality demonstrating particular robustness under topological enhancement ($5\times$ advantage over proxy strategies);
  • Mechanistic distinction: Centrality-based stimulation preserves temporal signal fidelity while proxy strategies amplify broadband noise.